### v1

@@ Abstract
The increasing complexity of application requirements and the dynamic nature of the Cloud-Edge Continuum present significant challenges for efficient resource management. These challenges stem from the ever-changing infrastructure—characterized by additions, removals, and reconfigurations of nodes and links—and the variability of application workloads. Traditional centralized approaches struggle to adapt to these changes due to their static nature, while decentralized solutions face challenges such as limited global visibility and coordination overhead. This paper proposes a hybrid decentralized framework for dynamic application placement and resource management. The framework utilizes Graph Neural Networks (GNNs) to embed resource and application states, enabling comprehensive representation and efficient decision-making. It employs a collaborative multi-agent reinforcement learning (MARL) approach, where local agents optimize resource management in their neighborhoods and a global orchestrator ensures system-wide coordination. By combining decentralized application placement with centralized oversight, our framework addresses the scalability, adaptability, and accuracy challenges inherent in the Cloud-Edge Continuum. This work contributes to the development of decentralized application placement strategies, the integration of GNN embeddings, and collaborative MARL systems, providing a foundation for efficient, adaptive and scalable resource management.
Index Terms:
AI, Continuum Computing, Dynamic Resource Management, Decentralized Application Placement, Multi-Agent Reinforcement Learning
@@ I Introduction
The Cloud-Edge Continuum has become a cornerstone of modern distributed computing, combining centralized cloud resources with edge and IoT devices to support diverse application requirements[1]. Applications in this Continuum range from real-time IoT services to computationally intensive data processing tasks, requiring resource management strategies that can balance latency, scalability, and efficiency[2]. However, the dynamic nature of the Continuum—marked by fluctuating workloads, evolving network topologies, and heterogeneous resource capabilities—introduces significant challenges for effective resource allocation[21].
Traditional resource management strategies typically rely on centralized architectures, which provide a global perspective but face challenges in adapting to dynamic changes due to high communication overhead and latency [16]. Decentralized approaches, while scalable and adaptive, face limitations such as inconsistent global state awareness and coordination challenges among agents[25]. Bridging these gaps necessitates a hybrid approach that integrates decentralized decision-making with a degree of centralized oversight to ensure both local efficiency and global optimization.
This paper proposes a decentralized application placement strategy for resource management in the Cloud-Edge Continuum. The framework leverages GNN to encode the states of resources and application components, providing a powerful representation for dynamic decision-making. Leveraging a collaborative MARL approach, local agents focus on optimizing resource management at the neighborhood level, while a global orchestrator facilitates coordination and high-level policy alignment. This integration of GNN-based embeddings and MARL enables the framework to balance scalability, adaptability, and accuracy in resource management. Below, we highlight the key contributions of this work:
  • Decentralized Application Placement Strategy: We propose a decentralized framework that allows local agents to make autonomous decisions while maintaining coordination with a global orchestrator.
  • Integration of GNN Embeddings: The use of GNNs enables accurate representation of resource network states and application requirements, enhancing decision-making at both local and global levels.
  • Collaborative MARL: By employing a MARL approach, we enable collaboration between local and global agents to achieve system-wide optimization while addressing scalability and adaptability challenges.
The rest of the paper is structured as follows: we begin by outlining the Problem Settings in section II, highlighting the key challenges and requirements of dynamic resource management in the Cloud-Edge Continuum. Next, in section III, we review existing approaches, including optimization-based methods, AI-driven techniques, decentralized strategies, and simulation tools for the Continuum. In section IV, we present our proposed framework, detailing the system model components including the application, resource, and cost models, along with the objective function. Section V describes the implementation, including local and global agents, their observation and action spaces, optimization techniques, and the training algorithm. Section VI explains the simulation environment, key evaluation metrics, and comparisons with centralized approaches. Finally, section VII outlines potential extensions, such as scaling to larger topologies, incorporating dynamic features, and validating in real-world scenarios.
@@ II Problem Settings
Traditional Cloud computing struggles to meet the quality-of-service (QoS) demands of IoT due to the scale, mobility, and geographical distribution of devices, as well as the need for low latency and localized processing[28]. The Cloud-Edge Continuum represents an evolution beyond traditional Cloud computing to address challenges arising from the rapid expansion of IoT devices and the vast amounts of data they generate. Unlike the cloud-centric model, the Continuum integrates a wide range of distributed computing resources across multiple layers, including cloud data centers, fog nodes, edge devices, and IoT endpoints[3]. This architecture aims to support diverse application requirements, including ultra-low latency for real-time interactions, high bandwidth for handling data-intensive tasks, distributed computing to enable parallel processing across multiple nodes, localized processing for data privacy and reduced network dependency, scalability to accommodate fluctuating workloads, and energy efficiency for sustainable operation in resource-constrained environments. However, the Continuum introduces several challenges, including the need for efficient resource discovery, monitoring, and management, as well as the coordination of application execution across heterogeneous resources and environments[5].
To evaluate resource management policies within the Cloud-Edge Continuum, three distinct models and a specified objective function are typically used:
  • Resource Model: represent the computational, storage, and network resources available across the Continuum, They capture device capabilities, network topology, resource heterogeneity, and dynamic changes such as variations in resource availability due to device mobility, workload fluctuations, or failures, providing a comprehensive view of the distributed environment.
  • Application Models: defines the structure and interactions within multi-services applications, including the topology of their components and communication paths. They capture the specific requirements of tasks or communication, such as latency sensitivity, bandwidth demands, and computational complexity. This ensures application constraints and performance metrics are accurately incorporated, enabling precise evaluation of how resource management policies impact application performance within the Continuum.
  • Cost Model: evaluate resource management expenses in the Continuum, considering computation and communication costs related to latency. Computation costs encompass the time required for processing tasks, while communication costs include the delay involved in transferring data, influenced by data size, bandwidth usage, and network conditions. These models provide a clear framework to balance performance and responsiveness in resource management policies.
  • Objective Function: focus on optimizing specific goals, such as minimizing latency, energy consumption, or operational costs while ensuring QoS and scalability. These models help identify trade-offs and balance competing objectives.
The dynamic nature of the Cloud-Edge Continuum, characterized by high variability, heterogeneity, and unpredictability, necessitates the use of autonomic control loop such as MAPE (Monitoring, Analysis, Planning, Execution) for effective resource management[19]. Models like resource, application, and cost models are integrated into the MAPE loop to guide decision-making in environments where resource states, and application demands fluctuate frequently.
@@ III-A Optimization based approaches
Directly solving the optimization problem through Mixed Integer Linear Programming (MILP) [13] is a well researched problem area that has theoretical guarantees to find the optimal solution, but is only suitable for small scale resource management problems. Heuristic-based algorithms instead pose the task as a multi-dimensional bin-packing problem such as firstfit [12], bestfit [4], worstfit[7] and round-robin placement[23] strategies which have long been used for resource allocation tasks. However, since the heuristics are only an approximation, these methods have no kind of optimality-guarantee. Both of these methods face the tradeoff between being compute heavy and providing non optimal answers, therefore they are only suitable for resource management on a small scale, or when strong constraints are placed on the problem structure.
III-B AI approaches
The dynamic and complex nature of the Cloud-Edge Continuum has sparked interest in leveraging AI approaches for resource management. As opposed to the aforementioned optimization-based approaches which often struggle to adapt to rapid changes in network topology and fluctuating application workloads, AI-based approaches such as DRL and GNNs have emerged as promising alternatives for their ability to approximate optimal solutions with reduced computational cost and enhanced adaptability.
DRL in Dynamic Resource Management: DRL has been widely explored for dynamic resource allocation, task scheduling and autoscaling in Cloud-Edge Continuum environments. In these settings, DRL agents learn policies by interacting with the environment, aiming to maximize cumulative rewards that are composed of performance metrics such as latency, throughput or cost. [17] proposed a DRL-based scheduler which adapts to workload variations in real-time, outperforming traditional heuristics-based methods. Whilst DRL-based methods excel with high-dimensional data, it can be beneficial to use compact representations of the data through embeddings.
When considering the centralized control approach, a single reinforcement learning agent with global view may often suffice, but in the instance of decentralized control, multi-agent reinforcement learning (MARL) must instead be considered. In [22], they use multi-agent DRL in wireless networks to enable simultaneous decisions to be made in a distributed manner. The framework is based on deep independent Q-Learning (IQL), but also accounts for a variable number of users and network densities whilst using a fixed-size neural network. However, a drawback of this approach is that agents learn in a non-stationary environments despite the effects of other agents actions on the environment and reward locally [14]. In their paper, they propose to avoid the issues of IQL through using a shared reward function to align agents and reduce learning instability.
To improve the representation of data for a DRL agent to be able to effectively learn at scale, embeddings such as those created with GNNs (e.g. Graph Convolutional Networks(GCNs)[32] and GraphSAGE[9]) can be used to manage variable size graphs as input. In [27], they leverage GNNs in a decentralized setting to better model the function of expected rewards in a larger DRL framework. A GNN can portray a very complex representation for a system comprising of connected nodes whilst allowing for a distributed implementation. However, a drawback of GNNs is the requirement to rebuild the network when a node is added or removed, thereby adding large computational and time cost. In our approach, we instead mask unused nodes in the DRL agent state space, thereby reducing the need to retrain the GNN.
III-C Decentralized Approaches
Traditionally in centralized approaches, there is a single global orchestrator with a comprehensive view of resources, making decisions to optimize resource allocation and performance across the Cloud-Edge Continuum[10]. In this setting, it is possible to achieve a globally optimal solution, however, this comes at the cost of scalability, increased latency and lack of resilience. Due to single controller, it can become a bottleneck for large scale systems which can slow the velocity at which decisions can be made.
In contrast, decentralized approaches distribute decision-making across multiple local agents, each with local information, that interact with neighboring nodes and agents. In [11], they use GNNs and MARL to enable decision making that incorporates local agent interactions. Each agent in the Vehicle-2-Anything (V2X) setting operates autonomously based on locally observed environment states, enabling efficient learning with reduced complexity. In this instance, they employ a weighted global reward to align local decisions to maximize their global objective function. However, the decentralization comes at a cost of no consistent view of the global state which can make lead to convergence at individual local states which may not align with the global optimum. Through leveraging GNNs and MARL pretraining at the local coordinator level along with the use of a global reward, we aim to offset the negative impacts of a fully decentralized approach.
III-D Dynamic management approaches
Dynamic resource management in the Cloud-Edge Continuum has garnered significant attention due to fluctuating application workloads and infrastructure dynamics. Research has explored adaptive resource allocation strategies to respond to changes in application demands, network conditions, and resource mobility. For instance, [29] introduced a dynamic resource provisioning algorithm that utilizes both edge and cloud resources with a latency-aware packet forwarding method for latency-sensitive applications. Many approaches integrate feedback mechanisms to monitor system states and enable real-time adjustments, optimizing resource utilization and reducing costs while meeting application-level QoS requirements. For example, [8] applies the MAPE control loop concept to design a resource provisioning framework for cloud environments. RL techniques have also been employed to address dynamism, such as the DRL-based dynamic resource management algorithm proposed by [6], which optimizes transmission power and computing resource allocation, effectively reducing task delay compared to conventional methods. Additionally, [24] employs invalid action masking to temporarily disable unavailable actions, effectively addressing the challenges of a variable action space. In this work, we also leverage RL and invalid action masking to address dynamic changes in the Cloud-Edge Continuum environment, but in a decentralized manner by employing localized multi-agent orchestration enabling adaptation to dynamic changes.
III-E Cloud-Edge Continuum Platform Simulation
Simulation frameworks play an essential role in evaluating novel strategies for resource management in the Cloud-Edge Continuum. This section discusses three simulators selected for their alignment with the specific scenario presented focusing on flexibility, scalability, and support for Edge-Cloud systems along with their extensibility for AI-based systems.
  • YAFS (Yet Another Fog Simulator)[15]: A modular simulator designed for edge and fog computing scenarios, supporting flexible resource modeling and dynamic application placement. YAFS excels at large-scale simulations in particular, enabling the evaluation of AI-based strategies in environments with high node density and dynamic workloads.
  • EdgeSimPy[30]: A lightweight Python-based simulator for edge computing environments, suitable for evaluating task placement and resource allocation strategies. Whilst less feature-rich than other simulators, it allows for rapid prototyping due to the integration with Python-native AI libraries.
  • ECLYPSE (Edge-CLoud pYthon Platform for Simulated runtime Environments) [18]: A simulation framework for distributed computing systems, with an emphasis on edge-cloud infrastructure and application behavior modeling. With a highly extensible framework, it offers detailed application behavior modeling and supports the interaction of Cloud and Edge components. ECLYPSE stands out for its balance of simplicity, extensibility and ability to simulate dynamic resource allocation strategies.
@@ IV Generic model
Our proposed model begins with a centralized use case, where GNNs are employed during the Analysis phase of the MAPE autonomic control loop. The GNN serves as a feature extractor, capturing the state of individual nodes while accounting for their logical inter-dependencies. These extracted features are then utilized by a centralized DRL agent, which has full observability of the entire system. The centralized agent focuses on optimizing global objectives, such as maximizing resource utilization and minimizing costs.
@@ IV-A System Model
The Cloud-Edge Continuum system comprises diverse resources distributed across multiple layers, including IoT devices, edge/fog nodes, and central clouds. These resources differ in computational power, memory, and network capabilities. Managing resources in this environment involves allocating application components to the available resources to optimize performance while meeting QoS requirements.
@@ IV-A1 Application Model
The application is modeled as a undirected graph $G_{A}=\langle C,E\rangle$, where:
  • $C$ denotes the set of vertices, expressed as $C=\{c_{1},c_{2},\dots,c_{N}\}$, where each vertex represents a component of the application. Following the microservice architecture, each component provides a unique service with specific requirements, including the number of CPU cores, GPU cores, amount of RAM, storage capacity, processing time for a specific device and the processing deadline. These requirements for each component $c_{i}$ are defined as $c_{i}=\{cpu_{i},gpu_{i},ram_{i},stor_{i},p_{i},ddl_{i}\}$.
  • $E$ represents the set of communication requirements$\{e_{1},e_{2},\dots,e_{N}\}$, defined as $e_{i}=\{lat(i,j),bw(i,j)\}$, between components. These requirements specify the communication link from $c_{i}$ to $c_{j}$. The interactions among services are further characterized by the maximum allowable message travel time (latency) and the minimum required message size for transmission.
@@ IV-A2 Resource Model
The resource model is defined as an undirected graph $G_{R}=\langle V,L\rangle$, where:
  • V = $\{v_{1},\dots,v_{M}\}$ where each item represents individual computational edge nodes or a central cloud data center. Each resource node $v_{i}=\{cpu_{i},gpu_{i},ram_{i},stor_{i},pt_{i},aval_{i}\}$ is characterized by its available capacity including resources such as CPU cores, GPU cores, amount of RAM, storage capacity, device response time, and its availability indicator.
  • L =$\{l_{i,j}\}$ represents the link between resource $v_{i}$ and resource $v_{j}$, with network capacity $l_{i,j}=\{lat(i,j),bw(i,j)\}$ comprised of packet delivery time along with the maximum bandwidth.
@@ IV-A3 Cost Model
The cost model comprises the following elements:
  • Completion Time: for a given application component, the upper bound is defined as the sum of computation and communication time:
    $CT(c_{i})=T_{comp}(c_{i},v_{k})+T_{comm}(c_{i},v_{k},c_{j},v_{m})$ (1)
    Where:
    $T\_comp(c_{i},v_{k})=pt_{v_{k}}+pt_{c_{i}}(v_{k})$
    $T\_comm(c_{i},v_{k},c_{j},v_{m})=lat_{i,j}(msgsize_{i,j},bw_{i,j})$
    The total completion time for the application is defined, assuming a sequential workflow, as:
    $CT_{app}=\sum_{c_{i}\in C}CT(c_{i})$ (2)
  • Resource Utilization: the average resource utilization across all nodes and edges is given by:
    $RU(t)=\frac{1}{|V|}\sum_{v_{i}\in V}\left(\frac{cpu_{used}(v_{i},t)}{cpu_{total}(v_{i})}+\frac{gpu_{used}(v_{i},t)}{gpu_{total}(v_{i})}+\dots\right)$ (3)
  • SLA Violation Rate: measures the percentage of application components that fail to reach their deadline:
[…]
@@ IV-A4 Objective Function
The objective of the resource management policy is to maximize resource utilization while minimizing completion time and SLA violation rate. This can be expressed as:
[…]
Where:
  • $RU$ is the average resource utilization.
  • $CT_{app}$ is the total application completion time.
  • $SVR$ is the SLA violation rate.
  • $\alpha,\beta,\gamma$ are weighting factors that determine the importance of each metric.
In future iterations of this model, the economic cost of deployment will be included as an additional optimization factor. This will allow comparisons between centralized and decentralized deployment policies, considering trade-offs such as infrastructure costs, operational expenses, and energy consumption.
@@ V-A Agents
The resource management framework incorporates two levels of agents: initially pretrained local agents and a global agent with aggregated local information, which are then jointly trained together to optimize application placement in the Cloud-Edge Continuum. These agents operate with the following characteristics:
@@ V-A1 Local Agents
Local agents act as coordinators for specific subgraph at the neighborhood level, managing multiple nodes within their designated areas. They have the following features:
  • Partial Observability: Local agents have visibility only over the nodes in their neighborhood and the local environment, enabling them to respond quickly to dynamic changes (see Sect.V-B).
  • Pretrained Behavior: Local agents are pretrained using local reward functions that emphasize efficiency within their region, such as maximizing local resource utilization, minimizing latency, and reducing SLA violations.
  • Local Reward Function: Each local agent optimizes a reward function defined as:
[…]
    Where:
    • $RU_{local_{i}}$: Local resource utilization.
    • $CT_{local_{i}}$: Completion time for tasks within the neighborhood.
[…]
    • $\alpha,\beta,\gamma$: Weight parameters for balancing reward terms.

### v2

@@ Abstract
In the Cloud-Edge Continuum, dynamic infrastructure change and variable workloads complicate efficient resource management. Centralized methods can struggle to adapt, whilst purely decentralized policies lack global oversight. This paper proposes a hybrid framework using Graph Neural Network (GNN) embeddings and collaborative multi-agent reinforcement learning (MARL). Local agents handle neighbourhood-level decisions, and a global orchestrator coordinates system-wide. This work contributes to decentralized application placement strategies with centralized oversight, GNN integration and collaborative MARL for efficient, adaptive and scalable resource management.
@@ I Introduction
Traditional Cloud computing struggles to meet the quality-of-service (QoS) demands of IoT due to the scale, mobility, and geographical distribution of devices, as well as the need for low latency and localized processing[15]. The Cloud-Edge Continuum represents an evolution beyond traditional Cloud computing to address challenges arising from the rapid expansion of IoT devices and the vast amounts of data they generate. Unlike the cloud-centric model, the Continuum integrates a wide range of distributed computing resources across multiple layers, including cloud data centers, fog nodes, edge devices, and IoT endpoints[1]. This architecture aims to support diverse application requirements, including distributed computing to enable parallel processing across multiple nodes, scalability to accommodate fluctuating workloads, and energy efficiency for sustainable operation in resource-constrained environments. However, the Continuum introduces several challenges, including the need for efficient resource discovery, monitoring, and management, as well as the coordination of application execution across heterogeneous resources and environments[2].
This paper proposes a decentralized application placement strategy for resource management in the Cloud-Edge Continuum. The framework leverages GNN to encode the states of resources and application components, providing a powerful representation for dynamic decision-making. Leveraging a collaborative MARL approach, local agents focus on optimizing resource management at the neighborhood level, while a global orchestrator facilitates coordination and high-level policy alignment. This integration of GNN-based embeddings and MARL enables the framework to balance scalability, adaptability, and accuracy in resource management.
@@ II Problem Settings
To evaluate resource management policies within the Cloud-Edge Continuum, three distinct models and a specified objective function, defined in IV-A, are typically used:
  • Resource Model: depicts available computational, storage and network links capturing resource heterogeneity and dynamic changes (e.g. device capabilities)
  • Application Models: defines the structure and interactions within multi-service applications, including the topology of their components and communication paths.
  • Cost Model: Considering completion time (CT) (computation and communication costs), resource utilization (RU) and SLA Violation Rate (SVR).
  • Objective Function: optimizing balance of cost model parameters, such as RU, SVR and CT.
The dynamic nature of the Cloud-Edge Continuum, characterized by high variability, heterogeneity, and unpredictability, necessitates the use of autonomic control loops such as MAPE (Monitoring, Analysis, Planning, Execution) for effective resource management[10]. Models like resource, application, and cost models are integrated into the MAPE loop to guide decision-making in environments where resource states, and application demands fluctuate frequently.
@@ III Related Work
Optimization-based approaches directly solving the optimization problem through methods such as Mixed Integer Linear Programming (MILP) [7] is a well-researched problem area that has theoretical guarantees to find the optimal solution, but is only suitable for small-scale resource management problems. Heuristic-based algorithms instead pose the task as a multi-dimensional bin-packing problem[19] such as first-fit, best-fit, worst-fit, and round-robin placement [13] strategies which have long been used for resource allocation tasks.
Deep Reinforcement Learning (DRL) has been extensively explored for dynamic resource allocation, task scheduling, and autoscaling in Cloud-Edge Continuum environments. While centralized control can leverage a single DRL agent with a global view, decentralized control requires multi-agent reinforcement learning (MARL), as demonstrated in [12], which used deep independent Q-learning (IQL) for distributed decisions in wireless networks. However, IQL struggles with non-stationary environments caused by other agents’ actions, a challenge addressed in [8] through shared reward functions to improve stability. To handle high-dimensional data effectively, embeddings created using GNNs (e.g., GCNs [18] and GraphSAGE [5]) allow DRL agents to process variable-size graphs, as shown in [14]. While GNNs excel in representing complex systems, they require costly rebuilding when nodes are added or removed. Our approach mitigates this by masking unused nodes in the DRL agent state space, eliminating the need for frequent retraining.
Research on dynamic resource management has explored adaptive strategies to address changing application demands, network conditions, and resource mobility. For example, [16] proposed a latency-aware algorithm that uses edge and cloud resources for latency-sensitive applications, while [4] applied the MAPE control loop for real-time resource provisioning in cloud environments. RL techniques, such as the DRL-based algorithm by [3], have also been used to optimize resource allocation and reduce task delays more effectively than traditional methods. Decentralized approaches distribute decision-making across local agents with limited information, interacting with neighboring nodes and agents. In [6], GNNs and MARL enable autonomous decision-making in the V2X setting, where agents learn efficiently from local observations and align decisions using a weighted global reward to maximize the global objective. However, decentralization lacks a consistent global view, potentially leading to suboptimal convergence. By leveraging GNNs, MARL pretraining at the local coordinator level, and a global reward, we aim to mitigate the downsides of full decentralization. Additionally, MARL enables our approach to adapt to dynamic system conditions and to make flexible decisions.
@@ IV Generic model
Our proposed model begins with a centralized use case, where GNNs are employed during the Analysis phase of the MAPE autonomic control loop. The GNN serves as a feature extractor, capturing the state of individual nodes while accounting for their logical inter-dependencies. These extracted features are then utilized by a centralized DRL agent, which has full observability of the entire system.
@@ IV-A System Model
In the Cloud-Edge Continuum, distributed resources differ in computational power, memory, and network capabilities. Managing resources in this environment involves allocating application components to the available resources to optimize performance while meeting QoS requirements.
@@ IV-A1 Application Model
An application is modeled as an undirected graph $G_{A}=\langle C,E\rangle$:
  • $C=\{c_{1},c_{2},\dots,c_{N}\}$ contains application components (microservices). Each component $c_{i}$ requires a certain number of CPU/GPU cores, memory, storage, and has a deadline.
  • $E$ specifies communication constraints, including latency and bandwidth needs between components.
@@ IV-A2 Resource Model
The resource model is modeled as $G_{R}=\langle V,L\rangle$, where:
  • V = $\{v_{1},\dots,v_{M}\}$ represents computing nodes (e.g. edge devices, cloud data centres), each with available CPU/GPU, RAM, storage, and an availability indicator.
  • $L$ describes network links with latency and bandwidth capacities.
@@ IV-A3 Cost Model
  • Completion Time (CT) of a component $c_{i}$ on a resource $v_{k}$ combines computation and communication times. Computation costs encompass the time required for processing tasks, while communication costs include the delay involved in transferring data, influenced by data size, bandwidth usage, and network conditions. Total application completion time is the sum of all components’ times, the upper bound is defined as the sum of computation and communication time.
  • Resource Utilization (RU) is the average resource utilization across all nodes and edges considering CPU, GPU, etc.
  • SLA Violation Rate (SVR) measures the fraction of components missing their deadlines.
@@ IV-A4 Objective Function
Maximizes RU while minimizing CT and SVR:
$\text{Objective: }\max\left(\alpha\cdot RU-\beta\cdot CT_{app}-\gamma\cdot\mathit{SVR}\right)$ (1)
Here, $\alpha,\beta,\gamma$ are weighting factors that determine the importance of RU, CT, and SVR, respectively. Future work will incorporate economic factors (e.g. operational costs, energy usage) to compare centralized and decentralized policies.
@@ V-A Agents
We employ a hierarchical design with local and global agents that collaborate to optimize application placement in the Cloud-Edge Continuum:
@@ V-A1 Local Agents
  • Partial Observability: Each local agent manages a partially observable neighbourhood subgraph(see Sect.V-B).
  • Pretrained Behavior: Pretrained using local reward to optimize efficiency within their subgraph.
  • Local Reward Function:
    $R_{local_{i}}=\alpha\cdot RU_{local_{i}}+\beta\cdot\frac{1}{CT_{local_{i}}}+\gamma\cdot\frac{1}{\mathit{SVR}_{local_{i}}}$ (2)
    balances RU, CT, and SVR for a given local agent.